OpenStudy (anonymous):

HELP PLEASE :MEDAL AND FAN Donna's math quiz required her to solve, if possible, the following algebraic equation: 3x + 2y = y She solved the equation this way: 5x = y Which statement explains what Donna did or should have done? Question 3 options: She simplified the equation correctly. She added the variables incorrectly, and the answer is incorrect She should have divided both sides by x, and the answer is incorrect. She should have multiplied the coefficients together, and the answer is incorrect.

OpenStudy (dogzcatz):

She simplified the equation correctly: I would of done 3x=y=0 She should have divided both sides by x, and the answer is incorrect: you don't nessasarily need to She should have multiplied the coefficients together, and the answer is incorrect: yes she should of

OpenStudy (anonymous):

That is actually incorrect^ In order to solve this, let's look at the problem itself: 3x + 2y = y So, what do we want as an answer? _x = _y, right? The first step would be to move the y's and x's on their own sides of the equation. So, you would subtract 2y from both sides, 3x + 2y = y -2y = -2y _________________ Leaving you with 3x = -y. Now let's look at Donna's answer: 5x = y She did not simplify the equation correctly, because her answer doesn't match ours. How did we solve it? We solved it by using addition/subtraction of the variables. Now, look back at the answer options. Donna did not need to divide anything by x nor did she need to multiply the coefficients. Therefore, there is only one remaining answer: she made a mistake in adding/subtracting her variables. Hope this helps and don't forget to click the 'best response' button :)